# Hierarchies of direct powers, ultrapowers and cumulative powers

### 16 Mar 2026 v1

@@ 4. Ultrapowers and the hierarchy
Notice the construction is consistent in that for any $x\in B$, $h_{x}$ is only defined once: if $r_{\alpha^{+}}+y\in B_{\alpha}$ for some $y\in B_{\alpha}$, then $r_{\alpha^{+}}+y-y=r_{\alpha^{+}}\in B_{\alpha}$; if $r_{\alpha^{+}}\cdot y\in B_{\alpha}$, then $r_{\alpha^{+}}\cdot y\cdot y^{-1}\in B_{\alpha}$; and if $r_{\alpha^{+}}^{-1}\in B_{\alpha}$, then $r_{\alpha^{+}}\in B_{\alpha}$. Therefore, whenever $r_{\alpha^{+}}\not\in B_{\alpha}$, every element $x\in B_{\alpha^{+}}\setminus B_{\alpha}$ is such that $h_{x}\not\in e_{\alpha}$, and therefore $h_{x}$ is not defined in $e_{\alpha}$, so that it may be defined in $e_{\alpha^{+}}$. By the construction, we may see $r_{i}\in B_{\beta}$ iff $h_{r_{i}}\in e_{\beta}$, and for every $r\in B$, $h_{r}\in e_{|B|}$. We note $e:B\to\mathcal{A}^{A}/\mathcal{U};x\mapsto[h_{x}]$ is an embedding. To see it is injective, let $r,r^{\prime}\in B$ and $r\neq r^{\prime}$. If $r,r^{\prime}\in A$, then $h_{r}=\overline{r}$ and $\overline{r^{\prime}}=h_{r^{\prime}}$, which means $\{j\in A\mid h_{r}(j)=h_{r^{\prime}}(j)\}=A\in\mathcal{U}$, and thus $e(r)=[h_{r}]\neq[h_{r^{\prime}}]=e(r^{\prime})$. Otherwise, without loss of generality, let $r\in B\setminus A$ and $r^{\prime}\in A$. We have either $r<r^{\prime}$ or $r<r^{\prime}$. In either case, by the definition of $h_{r^{\prime}}$, $\{j\in A\mid h_{r}(j)=\overline{r}(j)=r=h_{r^{\prime}}(j)\}\not\in\mathcal{U}$ and therefore $e(r)=[h_{r}]=[\overline{r}]\neq[h_{r^{\prime}}]=e(r^{\prime})$. Let now $r,r^{\prime}\in B\setminus A$. Once again, without loss of generality, let $r<r^{\prime}$. Since $\mathcal{A}$ is dense in $\mathcal{B}$ there is $m\in A$ such that $r<m<r^{\prime}$. That means $m\in L_{r^{\prime}}$ but $m\not\in L_{r}$. Since $h_{r^{\prime}}(A)$ is cofinal in $L_{r^{\prime}}$, there is $m^{\prime}\in h_{r^{\prime}}(A)$ such that $m<m^{\prime}$, so since $h_{r^{\prime}}(A)\subseteq L_{r^{\prime}}$, $m^{\prime}\in L_{r^{\prime}}$, which means $r<m<m^{\prime}<r^{\prime}$. Let $t\in A$ be such that $h_{r^{\prime}}(t)=m^{\prime}$. Since $h_{r^{\prime}}$ is order-preserving, for any $t^{\prime}>t$, $h_{r^{\prime}}(t^{\prime})>h_{r}(t^{\prime})$. Thus, we have $\{j\in A\mid h_{r}(j)<h_{r^{\prime}}(j)\}$ is a tail of $\mathcal{A}$, and therefore $\{j\in A\mid h_{r}(j)<h_{r^{\prime}}(j)\}\subseteq\{j\in A\mid h_{r}(j)\neq h_{r^{\prime}}(j)\}\in\mathcal{U}$. Hence, $e(r)=[h_{r}]\neq[h_{r^{\prime}}]=e(r^{\prime})$. Certainly, $e(0)=[0]$ and $e(1)=[1]$. Consider $r_{i},r_{j}\in B$. Then, $r_{i},r_{j}\in B_{max\{i,j\}}$, which means $h_{r_{i}},h_{r_{j}}\in e_{max\{i,j\}}$. By the definition of $e_{max\{i,j\}}$, we have $e(r_{i})+e(r_{j})=[h_{r_{i}}]+[h_{r_{j}}]=[h_{r_{i}}+_{1}h_{r_{j}}]=[h_{r_{i}+_{1}r_{j}}]=e(r_{i}+r_{j})$ (and likewise for multiplication). Let now $r_{i}<r_{j}$. Since $\mathcal{A}$ is dense in $\mathcal{B}$, that means there is $a\in L_{r_{j}}$ such that for all $b\in L_{r_{i}}$, $b<a$. Since $h_{r_{j}}(A)$ is cofinal in $L_{r_{j}}$, there is $t\in A$ such that $h_{r_{j}}(t)>a$, and since $h_{r_{j}}$ is order preserving, for any $t^{\prime}>t$, $h_{r_{j}}(t^{\prime})>h_{r_{j}}(t)$. Since $h_{r_{i}}(A)\subseteq L_{r_{i}}$, that means for any $t^{\prime}>t$, $h_{r_{j}}(t^{\prime})>h_{r_{i}}(t^{\prime})$. Therefore, $\{j\in A\mid h_{r_{i}}(j)<h_{r_{j}}(j)\}$ is a tail of $\mathcal{A}$, and therefore in $\mathcal{U}$, which means $e(r_{i})=[h_{r_{i}}]<_{\mathcal{U}}[h_{r_{j}}]=e(r_{j})$. ∎
Let $\mathbb{Q}^{*}$ be some construction of the hyperrationals, and $\infty$ and $\varepsilon$ be its infinite and infinitesimal elements, respectively. It is well known the quotient ring $(\mathbb{Q}^{*}\setminus\infty)/\varepsilon$ is isomorphic to $\mathbb{R}$. However, the above result provides a direct embedding of $\mathbb{R}$ into the construction of $\mathbb{Q}^{*}$ in the present framework:
@@
    $\mathbb{R}\preceq\overline{\mathbb{Q}}^{\overline{\mathbb{Q}}}/\mathcal{U}\cong\mathbb{F}_{1}(\overline{\mathbb{Q}})$. [^19] Given $I_{0}=\mathbb{Q}$ (respectively, $I_{0}=\overline{\mathbb{Q}}$), and a adequate choice of ultrafilter defining $\sim_{1}$.
That is, the hyperrational field $\mathbb{Q}^{\mathbb{Q}}/\mathcal{U}$, although not being real closed, contains an isomorphic copy of the reals; and the hyperalgebraic field $\overline{\mathbb{Q}}^{\overline{\mathbb{Q}}}/\mathcal{U}$ is a real closed non-Archimidean field extension of the reals. Notice, by the construction in Theorem 4.41, the relevant embeddings are not unique – as should be expected by the usual construction $(\mathbb{Q}^{*}\setminus\infty)/\varepsilon$.
@@ 5. An application to real closed fields
For an ordered set $\mathcal{A}$ and $f,g\in A^{A}$, say $f$ dominates $g$ if there is $j\in A$ such that for all $k>j$, $f(k)>g(k)$. Given an order $\langle A,<\rangle$ and $L,R\subseteq A$, we write $L<R$, as usual, to denote that every element of $L$ is lesser than every element of $R$. Given $a\in A$, for simplicity we write $a<L$ for $\{a\}<L$. Call $B\subseteq A^{A}$ unbounded if there is no $h\in A^{A}$ such that $h$ dominates every element of $B$. Following [18], define $\mathfrak{b}_{A}$ to be the size of the smallest unbounded subset of $A^{A}$. [12] has shown that $\aleph_{0}<\mathfrak{b}_{\omega}$. We adapt here that proof to the more general case:
@@
By the definition of $\mathfrak{b}_{\mathcal{A}}$, we may see that:
@@
A structure $\mathcal{A}$ is $\kappa$-saturated if for any $X\subseteq A$ with $X<\kappa$, $\mathcal{A}$ realizes any complete type over XX. [^20] A type over $X\subseteq A$ is a set $\Phi$ of formulas with $n$ free variables and parameters in $X$ such that any finite subset of $\Phi$ is satisfiable in $\mathcal{A}$. A type is realized by $\mathcal{A}$ if there is a $n$-tuple in $A^{n}$ satisfying $\Phi$, and complete if for any $\varphi$ with parameters in $X$, either $\varphi\in\Phi$ or $\neg\varphi\in\Phi$.
@@
For $\beta\in\mathbf{On}$, an ordered set $A$ is an $\eta_{\beta}$-set if for any $L,R\subseteq A$ with $|L|,|R|<\aleph_{\beta}$ and $L<R$ there is $a\in A$ such that $L<a<R$. Likewise, a $\eta_{\mathbf{On}}$-class is a proper class that is a $\eta_{\beta}$-class for every $\beta\in\mathbf{On}$. [^21] That is, such that for any sets $X,Y\subset\mathbf{No}$ such that $X<Y$ there is $a\in\mathbf{No}$ such that $X<a<Y$.
@@
Notice particularly that if $|A|=\aleph_{0}$, under the assumption of GCH, $|\Upsilon_{\lambda}(\mathcal{A})|=\aleph_{\lambda}$, and is therefore unique up to order isomorphism ([11], pp. 180–185). [^22] One might see that by induction, since $|A_{0}|=|A|=\aleph_{0}<\mathrm{cof}(A_{1})\leq|A^{A}|=2^{\aleph_{0}}=\aleph_{1}$.
@@
Given Theorem 4.25, the above results implies that if $\mathcal{A}$ is an $\eta_{0}$-set, $\{\Upsilon_{\alpha}(\mathcal{A})\}_{\alpha\in\mathbf{On}}$ is the family of index sets generating the cumulative power hierarchy, and each ultrafilter $\{\mathcal{U}_{\alpha}\}_{\alpha\in\mathbf{On}}$ defining $\sim_{\alpha^{+}}$ is a tails ultrafilter, then for each strongly inaccessible $\aleph_{\lambda}$, $\mathbb{F}_{\lambda}(\mathcal{A})$ is an $\eta_{\lambda}$-set, and $\mathbb{F}_{\mathbf{On}}(\mathcal{A})$ is an $\eta_{\mathbf{On}}$-class. However, we may eliminate the dependence on ultrapowers of that sort of construction by showing that, for any ultrapower hierarchy constructed by taking each successor level as an ultrapower having the preceding level as its index set – as constructed above –, there is a corresponding hierarchy of cumulative powers constructed by taking each preceding level as the index set of its successor such that, once quotiented by the equivalence relations $\sim$ defined by appropriate choices of ultrafilter, its structures are isomorphic to their ultrapower counterparts.
@@
Notice that, under a suitable assumption, the proof of the above result essentially goes both ways – that is, to show a ultrapower hierarchy isomorphic to the hierarchy of $\mathbb{F}_{\beta}(\mathcal{A})$ may be constructed.
@@
We conjecture that, in fact, such restriction on the ultrafilters is not necessary, and so that the similarities between both constructions is such that the quotiented $\mathcal{F}$-hierarchies and ultrapower hierarchies constructed by having the previous levels as the index sets of the successor levels are equivalent, so that any structure obtained by one of the constructions may be obtained by the other in a straightforward manner.
Now, given an ultrapower hierarchy, notice regardless of the choice of family of index sets and ultrafilters, by Łos’s theorem we immediately have:
@@
More specifically, if $\mathcal{A}$ is a real closed field, $|A|=\aleph_{\alpha}$, and for each $\mathcal{U}_{\beta}$ over $\Upsilon_{\beta}(\mathcal{A})$ is countably incomplete and $|\Upsilon_{\beta}(\mathcal{A})|^{+}$-good, then each $\Upsilon_{\beta^{+}}(\mathcal{A})$ – and its correspondent $\mathbb{F}_{\beta^{+}}(\mathcal{A})$ – is a $\aleph_{\alpha+\beta^{+}}$-saturated hyperreal field. Notice, however, that by Theorem 5.16 we may see there is a always a straightforward choice of ultrafilter over $\mathcal{F}_{\beta}(\mathcal{A})$ that makes $\mathbb{F}_{\beta^{+}}(\mathcal{A})$ a hyperreal field; and furthermore, that by Theorem 5.13 and Corollary 5.14, it is straightforward to make $\mathbb{F}_{\lambda}(\mathcal{A})$ into a $\aleph_{\lambda}$-saturated hyperreal field, and likewise, $\mathbb{F}_{\mathbf{On}}(\mathcal{A})$ into a maximally saturated real closed field without the need of countably incomplete good ultrafilters. In fact, consider the following notion.
@@
Introduced by Conway [5], the surreal field $\mathbf{No}$ is an universally embedding real closed field that is a proper class. Besides the operations of multiplication and addition, the operations of division and exponentiation may be further defined [5], [6]. The following results make the assumption of Global Choice, which we shall denote by writing its acronym GC.
@@
By Proposition 5.21, $\Upsilon_{\beta}(\mathcal{A})$ is a real-closed field. By Corollary 5.25, $\mathcal{A}_{\beta}$ is $\aleph_{\beta}$-saturate, so by Proposition 5.9, it is an $\eta_{\beta}$-set. By Proposition 5.30, $\Upsilon_{\beta}(\mathcal{A})$ is then a $\aleph_{\beta}$-universally embedding real-closed field. The argument is similar for the other cases. ∎

### 18 Mar 2026 v2

@@ 4.3. Embedding an ultrapower into a direct power
Notice the construction is consistent in that for any $x\in B$, $h_{x}$ is only defined once: if $r_{\alpha^{+}}+y\in B_{\alpha}$ for some $y\in B_{\alpha}$, then $r_{\alpha^{+}}+y-y=r_{\alpha^{+}}\in B_{\alpha}$; if $r_{\alpha^{+}}\cdot y\in B_{\alpha}$, then $r_{\alpha^{+}}\cdot y\cdot y^{-1}\in B_{\alpha}$; and if $r_{\alpha^{+}}^{-1}\in B_{\alpha}$, then $r_{\alpha^{+}}\in B_{\alpha}$. Therefore, whenever $r_{\alpha^{+}}\not\in B_{\alpha}$, every element $x\in B_{\alpha^{+}}\setminus B_{\alpha}$ is such that $h_{x}\not\in e_{\alpha}$, and therefore $h_{x}$ is not defined in $e_{\alpha}$, so that it may be defined in $e_{\alpha^{+}}$. By the construction, we may see $r_{i}\in B_{\beta}$ iff $h_{r_{i}}\in e_{\beta}$, and for every $r\in B$, $h_{r}\in e_{|B|}$. We note $e:B\to\mathcal{A}^{A}/\mathcal{U};x\mapsto[h_{x}]$ is an embedding. To see it is injective, let $r,r^{\prime}\in B$ and $r\neq r^{\prime}$. If $r,r^{\prime}\in A$, then $h_{r}=\overline{r}$ and $\overline{r^{\prime}}=h_{r^{\prime}}$, which means $\{j\in A\mid h_{r}(j)\neq h_{r^{\prime}}(j)\}=A\in\mathcal{U}$, and thus $e(r)=[h_{r}]\neq[h_{r^{\prime}}]=e(r^{\prime})$. Otherwise, without loss of generality, let $r\in B\setminus A$ and $r^{\prime}\in A$. We have either $r<r^{\prime}$ or $r>r^{\prime}$. In either case, by the definition of $h_{r^{\prime}}$, $\{j\in A\mid h_{r}(j)=\overline{r}(j)=r=h_{r^{\prime}}(j)\}\not\in\mathcal{U}$ and therefore $e(r)=[h_{r}]=[\overline{r}]\neq[h_{r^{\prime}}]=e(r^{\prime})$. Let now $r,r^{\prime}\in B\setminus A$. Once again, without loss of generality, let $r<r^{\prime}$. Since $\mathcal{A}$ is dense in $\mathcal{B}$ there is $m\in A$ such that $r<m<r^{\prime}$. That means $m\in L_{r^{\prime}}$ but $m\not\in L_{r}$. Since $h_{r^{\prime}}(A)$ is cofinal in $L_{r^{\prime}}$, there is $m^{\prime}\in h_{r^{\prime}}(A)$ such that $m<m^{\prime}$, so since $h_{r^{\prime}}(A)\subseteq L_{r^{\prime}}$, $m^{\prime}\in L_{r^{\prime}}$, which means $r<m<m^{\prime}<r^{\prime}$. Let $t\in A$ be such that $h_{r^{\prime}}(t)=m^{\prime}$. Since $h_{r^{\prime}}$ is order-preserving, for any $t^{\prime}>t$, $h_{r^{\prime}}(t^{\prime})>h_{r}(t^{\prime})$. Thus, we have $\{j\in A\mid h_{r}(j)<h_{r^{\prime}}(j)\}$ is a tail of $\mathcal{A}$, and therefore $\{j\in A\mid h_{r}(j)<h_{r^{\prime}}(j)\}\subseteq\{j\in A\mid h_{r}(j)\neq h_{r^{\prime}}(j)\}\in\mathcal{U}$. Hence, $e(r)=[h_{r}]\neq[h_{r^{\prime}}]=e(r^{\prime})$. Certainly, $e(0)=[0]$ and $e(1)=[1]$. Consider $r_{i},r_{j}\in B$. Then, $r_{i},r_{j}\in B_{max\{i,j\}}$, which means $h_{r_{i}},h_{r_{j}}\in e_{max\{i,j\}}$. By the definition of $e_{max\{i,j\}}$, we have $e(r_{i})+e(r_{j})=[h_{r_{i}}]+[h_{r_{j}}]=[h_{r_{i}}+_{1}h_{r_{j}}]=[h_{r_{i}+_{1}r_{j}}]=e(r_{i}+r_{j})$ (and likewise for multiplication). Let now $r_{i}<r_{j}$. Since $\mathcal{A}$ is dense in $\mathcal{B}$, that means there is $a\in L_{r_{j}}$ such that for all $b\in L_{r_{i}}$, $b<a$. Since $h_{r_{j}}(A)$ is cofinal in $L_{r_{j}}$, there is $t\in A$ such that $h_{r_{j}}(t)>a$, and since $h_{r_{j}}$ is order preserving, for any $t^{\prime}>t$, $h_{r_{j}}(t^{\prime})>h_{r_{j}}(t)$. Since $h_{r_{i}}(A)\subseteq L_{r_{i}}$, that means for any $t^{\prime}>t$, $h_{r_{j}}(t^{\prime})>h_{r_{i}}(t^{\prime})$. Therefore, $\{j\in A\mid h_{r_{i}}(j)<h_{r_{j}}(j)\}$ is a tail of $\mathcal{A}$, and therefore in $\mathcal{U}$, which means $e(r_{i})=[h_{r_{i}}]<_{\mathcal{U}}[h_{r_{j}}]=e(r_{j})$. ∎
@@
    $\mathbb{R}\preceq\overline{\mathbb{Q}}^{\overline{\mathbb{Q}}}/\mathcal{U}\cong\mathbb{F}_{1}(\overline{\mathbb{Q}})$. [^20] Given $I_{0}=\mathbb{Q}$ (respectively, $I_{0}=\overline{\mathbb{Q}}$), and an adequate choice of ultrafilter over it defining $\sim_{1}$.
That is, the hyperrational field $\mathbb{Q}^{\mathbb{Q}}/\mathcal{U}$, although not being real closed, contains an isomorphic copy of the reals; and the hyperalgebraic field $\overline{\mathbb{Q}}^{\overline{\mathbb{Q}}}/\mathcal{U}$ is a real closed non-Archimedean field extension of the reals. Notice, by the construction in Theorem 4.41, the relevant embeddings are not unique – as should be expected by the usual construction $(\mathbb{Q}^{*}\setminus\infty)/\varepsilon$.
@@ 5. An application to real closed fields
For an ordered set $\mathcal{A}$ and $f,g\in A^{A}$, say $f$ dominates $g$ if there is $j\in A$ such that for all $k>j$, $f(k)>g(k)$. Given an order $\langle A,<\rangle$ and $L,R\subseteq A$, we write $L<R$, as usual, to denote that every element of $L$ is less than every element of $R$. Given $a\in A$, for simplicity we write $a<L$ for $\{a\}<L$. Call $B\subseteq A^{A}$ unbounded if there is no $h\in A^{A}$ such that $h$ dominates every element of $B$. Following [19], define $\mathfrak{b}_{A}$ to be the size of the smallest unbounded subset of $A^{A}$. [11] has shown that $\aleph_{0}<\mathfrak{b}_{\omega}$. We adapt here that proof to the more general case:
@@
By the definition of $\mathfrak{b}_{\mathcal{A}}$, we may immediately see that:
@@
A structure $\mathcal{A}$ is $\kappa$-saturated if for any $X\subseteq A$ with $|X|<\kappa$, $\mathcal{A}$ realizes any complete type over XX. [^21] A type over $X\subseteq A$ is a set $\Phi$ of formulas with $n$ free variables and parameters in $X$ such that any finite subset of $\Phi$ is satisfiable in $\mathcal{A}$. A type is realized by $\mathcal{A}$ if there is an $n$-tuple in $A^{n}$ satisfying $\Phi$, and complete if for any $\varphi$ with parameters in $X$, either $\varphi\in\Phi$ or $\neg\varphi\in\Phi$.
@@
For $\beta\in\mathbf{On}$, an ordered set $A$ is an $\eta_{\beta}$-set if for any $L,R\subseteq A$ with $|L|,|R|<\aleph_{\beta}$ and $L<R$ there is $a\in A$ such that $L<a<R$. Likewise, $A$ is an $\eta_{\mathbf{On}}$-class ig it is a proper class that is a $\eta_{\beta}$-class for every $\beta\in\mathbf{On}$. [^22] That is, such that for any sets $X,Y\subset A$ such that $X<Y$ there is $a\in A$ such that $X<a<Y$.
@@
Notice particularly that if $|A|=\aleph_{0}$, under the assumption of GCH, $|\Upsilon_{\lambda}(\mathcal{A})|=\aleph_{\lambda}$, and is therefore unique up to order isomorphism ([10], pp. 180–185). [^23] One might see, since we have GCH, that by induction, $|A_{0}|=|A|=\aleph_{0}<\mathrm{cof}(A_{1})\leq|A^{A}|=2^{\aleph_{0}}=\aleph_{1}$.
@@
Given Theorem 4.25, the above results imply that if $\mathcal{A}$ is an $\eta_{0}$-set, $\{\Upsilon_{\alpha}(\mathcal{A})\}_{\alpha\in\mathbf{On}}$ is the family of index sets generating the cumulative power hierarchy, and each ultrafilter $\{\mathcal{U}_{\alpha}\}_{\alpha\in\mathbf{On}}$ defining $\sim_{\alpha^{+}}$ is a tails ultrafilter, then for each strongly inaccessible $\aleph_{\lambda}$, $\mathbb{F}_{\lambda}(\mathcal{A})$ is an $\eta_{\lambda}$-set, and $\mathbb{F}_{\mathbf{On}}(\mathcal{A})$ is an $\eta_{\mathbf{On}}$-class. However, we may eliminate the dependence on ultrapowers of that sort of construction by showing that, for any ultrapower hierarchy constructed by taking each successor level as an ultrapower having the preceding level as its index set – as constructed above –, there is a corresponding hierarchy of cumulative powers constructed by taking each preceding level as the index set of its successor such that, once quotiented by the equivalence relations $\sim$ defined by appropriate choices of ultrafilter, its structures are isomorphic to their ultrapower counterparts.
@@
Notice that, under a suitable assumption, the proof of the above result essentially goes both ways – that is, to show an ultrapower hierarchy isomorphic to the hierarchy of $\mathbb{F}_{\beta}(\mathcal{A})$ may be constructed.
@@
We conjecture that, in fact, such a restriction on the ultrafilters is not necessary, and so that the similarities between both constructions are such that the quotiented $\mathcal{F}$-hierarchies and ultrapower hierarchies constructed by having the previous levels as the index sets of the successor levels are equivalent, so that any structure obtained by one of the constructions may be obtained by the other in a straightforward manner.
Now, given an ultrapower hierarchy, notice regardless of the choice of family of index sets and ultrafilters, by Łoś’s theorem we immediately have:
@@
More specifically, if $\mathcal{A}$ is a real closed field, $|A|=\aleph_{\alpha}$, and for each $\mathcal{U}_{\beta}$ over $\Upsilon_{\beta}(\mathcal{A})$ is countably incomplete and $|\Upsilon_{\beta}(\mathcal{A})|^{+}$-good, then each $\Upsilon_{\beta^{+}}(\mathcal{A})$ – and its corresponding $\mathbb{F}_{\beta^{+}}(\mathcal{A})$ – is a $\aleph_{\alpha+\beta^{+}}$-saturated hyperreal field. Notice, however, that by Theorem 5.16 we may see there is always a straightforward choice of ultrafilter over $\mathcal{F}_{\beta}(\mathcal{A})$ that makes $\mathbb{F}_{\beta^{+}}(\mathcal{A})$ a hyperreal field; and furthermore, that by Theorem 5.13 and Corollary 5.14, it is straightforward to make $\mathbb{F}_{\lambda}(\mathcal{A})$ into a $\aleph_{\lambda}$-saturated hyperreal field, and likewise, $\mathbb{F}_{\mathbf{On}}(\mathcal{A})$ into a maximally saturated real closed field without the need for countably incomplete good ultrafilters. In fact, consider the following notion.
@@
Introduced by Conway [5], the surreal field $\mathbf{No}$ is a universally embedding real closed field that is a proper class. Besides the operations of multiplication and addition, the operations of division and exponentiation may be further defined [5], [18]. The following results make the assumption of Global Choice, which we shall denote by writing its acronym GC.
@@
By Proposition 5.21, $\Upsilon_{\beta}(\mathcal{A})$ is a real-closed field. By Corollary 5.25, $\Upsilon_{\beta}(\mathcal{A})$ is $\aleph_{\beta}$-saturated, so by Proposition 5.9, it is an $\eta_{\beta}$-set. By Proposition 5.30, $\Upsilon_{\beta}(\mathcal{A})$ is then an $\aleph_{\beta}$-universally embedding real-closed field. The argument is similar for the other cases. ∎